# Regurgitative Training: The Value of Real Data in Training Large Language Models

### v1

@@ 1 Introduction
Large language models (LLMs) are trained on inexplicably large amounts of data. Although the exact training datasets are undisclosed, popular LLMs such as ChatGPT, Llama, Claude, and Mistral are believed to have been trained on a combination of content on public Internet (e.g., Common Crawl), proprietary datasets licensed from third-parties, as well as crowd-generated data (Brown et al.,, 2020; Ouyang et al.,, 2022; Achiam et al.,, 2023; Touvron et al.,, 2023; Anthropic,, 2024). With their explosive successes come explosive adoption – people use LLMs in an ever increasing set of tasks, including writing (Noy and Zhang,, 2023; Chen and Chan,, 2023), coding (Chen et al.,, 2021), knowledge management (Lewis et al.,, 2020), scientific discovery (Bran et al.,, 2023; Vert,, 2023), and many more.
@@
A natural consequence of such pervasive use is that, going forward, a substantial amount of content online will be created (at least partially) by LLMs. When building the next-generation LLMs, data generated by existing LLMs is likely to enter the training datasets. This produces a scenario which we refer to as Regurgitative Training, where a new LLM is trained using data that are at least partially generated by itself or other LLMs. The overarching question we seek to answer in this paper is: how does regurgitative training affect the performance of LLMs?
@@
Regurgitative training may be inevitable. Indeed, there is evidence suggesting that a large part of the open web is already generated by machine translation models (LLMs included, Thompson et al.,, 2024). Even data that are supposed to be human-generated (e.g., manual labels on crowdsourcing platforms) are often generated by LLMs (Veselovsky et al.,, 2023). As LLMs get better, it will be increasingly hard to distinguish between LLM-generated data from human-generated data post-hoc (Yang et al.,, 2023). Addition, some LLM developers have explicitly chosen to inject LLM-generated data into their training datasets. For example, Apple acknowledges that its multi-modal LLM named MM1 has been trained on instruction-response pairs generated from GPT-4 and Llama (McKinzie et al.,, 2024).
@@
A priori, the impact of regurgitative training on LLM performance is unclear. On one hand, it represents an appealing opportunity to obtain large quantities of synthetic training data at relatively low costs, thereby offering a data quantity advantage. On the other hand, however, LLM-generated data may have lower quality than real, human-generated data – they may contain more errors (Shumailov et al.,, 2023) or suffer more from the “hallucination problem” (Rawte et al.,, 2023). [^1] Throughout the paper, we use “real data” or ”real human-generated data”, in contrast with synthetic LLM-generated data, to refer to data that are generated by an organic process (typically by humans). Importantly, we do not assume real, human-generated data to be completely error-free; instead, we are interested in the comparison of real vs. synthetic data in model training. In fact, although major players in the LLM arena, including Microsoft, Google and Meta, are all reported to use synthetic data for LLM training, the practice has caused doubts from mainstream media. [^2] Sources: Microsoft, Google and Meta Bet on Fake Data to Build AI Models, The AI Revolution Is Already Losing Steam, For Data-Guzzling AI Companies, the Internet Is Too Small, and AI-Generated Data Can Poison Future AI Models.
@@
Across different generative tasks and model settings, we consistently observe that LLMs with regurgitative training underperforms those trained with real data. Given the same base model, training with more real data typically improves performance, whereas training with more LLM-generated data leads to quickly plateaued performance or even performance drops. Moreover, training with even a small proportion of real data is enough to outperform training only with LLM-generated data. The performance disadvantages of regurgitative training are especially pronounced when the LLM responsible for generating training data is not good at the task.
We also perform several textual analyses to make sense of regurgitative training’s performance disadvantages. As can be expected, errors in LLM-generated data is one of the culprits. Interestingly, we find evidence that it may not be the only contributing mechanism. Specifically, LLM-generated data exhibit lower degrees of lexical diversity than real data, echoing several recent research (Padmakumar and He,, 2023; Doshi and Hauser,, 2023; Anderson et al.,, 2024) in the contexts of academic writing or creative ideation. Lower lexical diversity in LLM-generated data may also partially explain the performance disadvantages of regurgitative training.
In light of these findings, we propose and evaluate three different strategies to more carefully leverage LLM-generated data in regurgitative training. The first strategy borrows from the semi-supervised learning literature and prioritizes using high-quality LLM-generated data over low-quality ones, where “quality” is gauged either by prediction confidence or by an external supervised learning model. Second, as an attempt to address the diversity deficit of LLM-generate data, we combine data generated by a mixture of different LLMs in training. Finally, the third strategy makes use of “AI detectors”, i.e., classification models that try to distinguish between LLM- vs. human-generated content. We train and deploy a capable AI detector model on LLM-generated data, and use LLM-generated data in regurgitative training in the order of predicted probability of being generated by humans (i.e., prioritizing LLM-generated data that resemble human-generated data). Our results demonstrate that all three strategies have some power to improve the performance of regurgitative training. In a few cases with transformer models trained from scratch, the improved performance even surpasses training with real data. Meanwhile, effectiveness of these strategies tends to be small in the fine-tuning setting, and none of them can fully close the gap from training with real data. This further highlights the unique value of real, human-generated data in LLM training.
@@ 2 Relevant Literature
Our work is closely related to self training in the semi-supervised learning literature and data augmentation in the deep learning literature, both of which are briefly reviewed in this section. As will be discussed later, although regurgitative training in LLMs is fundamentally different from the conventional schemes of self-training or data augmentation, both offer some valuable ideas that can inform our understanding of regurgitative training as well as potential approaches to manage its performance downsides.
@@ 2.1 Self Training
There is an extensive literature on self training, both in traditional machine learning (see Pise and Kulkarni,, 2008; Triguero et al.,, 2015, for surveys of the topic) and in modern deep learning (e.g., Xie et al., 2020b, ). One key insights from this body of work is that the effectiveness of self training depends heavily on the ability to accurately estimate “prediction confidence”. Because predictions with high confidence are used as “pseudo-labels”, having accurate confidence scores imply that the “pseudo-labels” are more likely to be correct (i.e., the same as ground-truth labels).
Regurgitative training of LLMs resembles self training in that model-predicted pseudo labels are used to further train the model. However, it is unclear whether the conventional wisdom of self training would still apply for in the case of regurgitative training, because of the difficulty in assessing confidence of LLM outputs (Lin et al.,, 2023). In classification tasks, predicted class probabilities naturally serve as the measure to quantify the uncertainty in a classifier’s predictions. However, LLMs are much more complex than classifiers – they generate multi-token answers in response to prompts. Unless in highly restricted scenarios (e.g., evaluating a single-digit response to the question “what is 2+2”), it is generally not straightforward to define or measure confidence in LLM outputs. As a result, current LLMs do not automatically produce confidence scores for their responses and uncertainty quantification in LLMs remain an open question with many ongoing research, including asking LLMs themselves for confirmation (“self-reflection”, Chen and Mueller,, 2023), re-running the same prompt multiple times and measure internal consistency among responses (Kotelanski et al.,, 2023), and tapping into human expertise (Shankar et al.,, 2024).
@@ 2.2 Data Augmentation
Data augmentation represents another approach to enrich potentially limited labeled data. It works by injecting noises into existing labeled data instances to artificially create new data instances that can be assigned the same labels. In language tasks, a common data augmentation strategy is back-translation (Yu et al.,, 2018), where a sentence is first translated into a different language then back to the original language to achieve paraphrasing. In vision tasks, data augmentation may involve image transformation techniques such as rotation, color / contrast modification, etc. (Cubuk et al.,, 2020; Xie et al., 2020a, ). These augmentation strategies can benefit model performance if the injected noises do not change the labels, thereby creating more training data with valid labels.
@@ 2.3 Regurgitative Training
Regurgitative training of generative AI models represents a new problem that has only begun to receive scholarly attention very recently. The earliest work we could identify is Shumailov et al., (2023), which document that using model-generated data to train next-generation models can create irreversible performance losses, a phenomenon they term “model collapse”. They demonstrate this in common generative AI architectures such as variational autoencoders, Gaussian mixture models, and small-scale LLMs. Moreover, they provide theoretical intuitions that model collapse arises due to errors in model-generated data, which accumulates over more iterations of regurgitative training. Subsequently, the phenomenon of model collapse has also been observed in generative image models (Alemohammad et al.,, 2023; Bertrand et al.,, 2023).
In the meantime, efforts to mitigate model collapse are underway. Bertrand et al., (2023) show that model collapse can be avoided if (i) the proportion of real data is sufficiently high and (ii) model-generated data approximate the distributions of real data well enough. Furthermore, Gerstgrasser et al., (2024) propose to alleviate model collapse by “accumulating data”; that is, use the totality of real and model-generated data (rather than just the model-generated data) to train new models.
We build upon this nascent stream of research and aim to make several distinct contributions. First, we consider regurgitative training of a LLM not only by data generated by itself, but also by other LLMs with varying degrees of capabilities. This is already taking place in practice (e.g., McKinzie et al.,, 2024) but has not been systematically explored in the literature. Second, prior work such as Shumailov et al., (2023) focused on early versions of generative models (e.g., non-transformer-based models or small pre-trained models). Instead, we carry out comprehensive experiments with leading LLMs at the time of research (e.g., GPT-4 and Llama2) as well as transformer models trained from scratch, thereby providing a more up-to-date understanding of regurgitative training. Finally, we propose several new mitigation strategies beyond what has been tested so far, and empirically evaluate their effectiveness.
@@ 3 Performance Impact of Regurgitative Training
In this section, we aim to understand how regurgitative training affect the performance of an LLM through two sets of experiments, respectively constructed to reflect two representative practices in LLM training: (i) fine-tuning and (ii) training from scratch. Fine-tuning allows users to adapt an existing LLM to their own use cases and, as mentioned before, is a widely adopted practice in the industry (e.g., McKinzie et al.,, 2024). We expect a lot of LLM training will take the form of fine-tuning, because training a state-of-the-art LLM from scratch is highly complex and resource-intensive. Meanwhile, we also consider the case of training a smaller-scale LLM from scratch, which may be necessary for companies that cannot leverage third-party LLMs due to data security and privacy issues.
@@ 3.1 Experiments with Fine-Tuning
In the first term, $c$ is the length of hypothesis translation and $r$ is the “effective” length of reference translations (defined as the length of the one reference translation that best matches the hypothesis translation). This term serves as a “brevity penalty” that assigns a higher score for a better match in lengths between hypothesis and reference translations. In the second term, $p_{n}$ denotes the $n$-gram precision and is defined as
$p_{n}=\frac{\sum_{n\mbox{-}gram\in hypothesis}Count_{matched}(n\mbox{-}gram)}{\sum_{ngram\in hypothesis}Count(n\mbox{-}gram)}$ (2)
where $Count_{matched}(n\mbox{-}gram)$ counts the number of $n$-gram matches between the hypothesis translation and reference translations. In Equation (1), the $n$-gram precision scores are then weighted by $w_{n}$ (e.g., uniform weighting $w_{n}=\frac{1}{N}$) to compute the overall BLEU score.
@@
We show the results in Figure 1. Each line represents the translation perform of a particular model over five fine-tuned batches. The $X$-axis indicates batch index, marking the number of data instances utilized in the fine-tuning process. The $Y$-axis represents the BLEU score, where a higher value corresponds to better translation performance.
@@
From the figure, it is evident that the performance of fine-tuning with LLM-generated data (both from the baseline LLM itself and from other LLMs) clearly lags behind the performance of fine-tuning with real human-generated data. Moreover, regurgitative training with different LLMs have differential performance impact. Fine-tuning with data generated by GPT-3.5 itself does not significantly change performance, and fine-tuning with GPT-4 generated data only results in barely noticeable performance improvement over the baseline model (i.e., at point 0 on the $X$-axis). However, fine-tuning with LLAMA2 generated data significantly degrades performance compared to the baseline. This is likely because the three LLMs have different translation capabilities. Since we have the ground-truths translations for the 5,000 fine-tuning data, we can directly compute the BLEU scores of translations generated by the three LLMs, and indeed find GPT-4 to be the best ($BLEU=0.3454$), followed by GPT-3.5 ($BLEU=0.3428$) and LLAMA2 ($BLEU=0.2417$).
These results underscore the overall negative, and potentially detrimental, effects of regurgitative training. Compared to training with real data, regurgitative training largely stalls learning. Regurgitative training with a better-performing LLM improves performance only marginally and is not sufficient to catch up with the performance on real data. Worse yet, regurgitative training with a less capable LLM can significantly hurt performance.
Note that in the above experiments, we use utilize a small set of data for fine-tuning. This decision stems from the remarkable few-shot learning capabilities of modern LLMs (Brown et al.,, 2020). In addition, we also conduct a robustness check to understand whether the performance of regurgitative training may be different if more fine-tuning data are available. Specifically, we augment the fine-tuning data size by 20 times, to a total of 100,000, and incrementally add 10,000 per batch. For efficiency and cost considerations, we only run fine-tuning with real data and data generated by GPT-3.5 itself. We then evaluate each fine-tuned models on the same testing data as before. The results are presented in Figure 2. We again observe that regurgitative training is unable to improve translation performance and substantially underperforms training with real data.

### v2

@@ 1 Introduction
Large language models (LLMs) are trained on inexplicably large amounts of data. Although the exact training datasets are undisclosed, popular LLMs such as ChatGPT, LLAMA, Claude, and Mistral are believed to have been trained on a combination of content on public Internet (e.g., Common Crawl), proprietary datasets licensed from third-parties, as well as crowd-generated data (Brown et al.,, 2020; Ouyang et al.,, 2022; Achiam et al.,, 2023; Touvron et al.,, 2023; Anthropic,, 2024). With their explosive successes come widespread adoption – people use LLMs in an ever increasing set of tasks, including writing (Noy and Zhang,, 2023; Chen and Chan,, 2023), coding (Chen et al.,, 2021), knowledge management (Lewis et al.,, 2020), scientific discovery (Bran et al.,, 2023; Vert,, 2023), and many more.
@@
A natural consequence of such pervasive use is that, going forward, a substantial amount of content online will be created (at least partially) by LLMs. When building the next-generation LLMs, data generated by existing LLMs are likely to enter the training datasets. This produces a scenario which we refer to as Regurgitative Training, where a new LLM is trained using data that are at least partially generated by itself or other LLMs. The overarching question we seek to answer in this paper is: how does regurgitative training affect the performance of LLMs?
@@
Regurgitative training may be inevitable. Indeed, there is evidence suggesting that a large part of the open web is already generated by machine translation models (LLMs included, Thompson et al.,, 2024). Even data that are supposed to be human-generated (e.g., manual labels on crowdsourcing platforms) are often generated by LLMs (Veselovsky et al.,, 2023). As LLMs get better, it will be increasingly hard to distinguish between LLM-generated data from human-generated data post-hoc (Yang et al.,, 2023). Additionally, some LLM developers have explicitly chosen to inject LLM-generated data into their training datasets, and empirically examine whether doing so can boost performance. For example, Apple acknowledges that its multi-modal LLM named MM1 has been trained on instruction-response pairs generated from GPT-4 and LLAMA(McKinzie et al.,, 2024).
@@
A priori, the impact of regurgitative training on LLM performance is unclear. On one hand, it represents an appealing opportunity to obtain large quantities of synthetic training data at relatively low costs, thereby offering a data quantity advantage. On the other hand, however, LLM-generated data may have lower quality than real, human-generated data – they may contain more errors (Shumailov et al.,, 2023) or suffer more from the “hallucination problem” (Rawte et al.,, 2023). [^1] Throughout the paper, we use “real data” or ”real human-generated data”, in contrast with synthetic LLM-generated data, to refer to data that are generated by an organic process (typically by humans). Importantly, we do not assume real, human-generated data to be completely error-free; instead, we are interested in the comparison of real vs. synthetic data in model training. Overall, the performance impact of regurgitative training is jointly affected by both quantity and quality of synthetic data. In fact, although major players in the LLM arena, including Microsoft, Google and Meta, are all reported to use synthetic data for LLM training, the practice has garnered doubts from mainstream media. [^2] Sources: Microsoft, Google and Meta Bet on Fake Data to Build AI Models, The AI Revolution Is Already Losing Steam, For Data-Guzzling AI Companies, the Internet Is Too Small, and AI-Generated Data Can Poison Future AI Models.
@@
Across different generative tasks and model settings, we consistently observe that LLMs with regurgitative training underperform those trained with real data. Given the same base model, training with more real data typically improves performance, whereas training with more LLM-generated data leads to quickly plateaued performance or even performance drops. In other words, the presumed performance advantage due to accessing large quantities of synthetic data is unrealized. Moreover, training with even a small proportion of real data is enough to outperform training only with LLM-generated data. The performance disadvantages of regurgitative training are especially pronounced when the LLM responsible for generating training data is not good at the task.
We also perform several textual analyses to make sense of regurgitative training’s performance disadvantages. As can be expected, errors in LLM-generated data is one of the culprits. Interestingly, we find evidence that it may not be the only contributing mechanism. Specifically, LLM-generated data exhibit lower degrees of lexical diversity than real data, echoing several recent research (Padmakumar and He,, 2023; Doshi and Hauser,, 2023; Anderson et al.,, 2024) in the contexts of academic writing or creative ideation. Lower lexical diversity in LLM-generated data also partially explains the performance disadvantages of regurgitative training.
In light of these findings, we propose and evaluate three different strategies to more carefully leverage LLM-generated data in regurgitative training. The first strategy borrows from the semi-supervised learning literature and prioritizes using high-quality LLM-generated data over low-quality data, where “quality” is gauged either by prediction confidence or by an external supervised learning model. Second, as an attempt to address the diversity deficit of LLM-generated data, we combine data generated by a mixture of different LLMs in training. Finally, the third strategy makes use of “AI detectors”, i.e., classification models that try to distinguish between LLM- vs. human-generated content. We train and deploy a capable AI detector model on LLM-generated data, and use LLM-generated data in regurgitative training in the order of predicted probability of being generated by humans (i.e., prioritizing LLM-generated data that resemble human-generated data). Our results demonstrate that all three strategies have some power to improve the performance of regurgitative training. In a few cases with transformer models trained from scratch, the performance improvements are quite significant. Meanwhile, effectiveness of these strategies tends to be small in the fine-tuning setting, and none of them can fully close the gap from training with real data. This further highlights the unique value of real, human-generated data in LLM training.
@@ 2 Relevant Literature
Our work is closely related to self training in the semi-supervised learning literature and data augmentation in the deep learning literature, both of which are briefly reviewed in this section. As will be discussed later, although regurgitative training in LLMs is fundamentally different from the conventional schemes of self training or data augmentation, both offer some valuable ideas that can inform our understanding of regurgitative training as well as potential approaches to manage its performance downsides.
@@ 2.1 Self Training
There is an extensive literature on self training, both in traditional machine learning (see Pise and Kulkarni,, 2008; Triguero et al.,, 2015, for surveys of the topic) and in modern deep learning (e.g., Xie et al., 2020b, ). One key insight from this body of work is that the effectiveness of self training depends heavily on the ability to accurately estimate “prediction confidence”. Because predictions with high confidence are used as “pseudo-labels”, having accurate confidence scores imply that the “pseudo-labels” are more likely to be correct (i.e., the same as ground-truth labels).
Regurgitative training of LLMs resembles self training in that model-predicted pseudo labels are used to further train the model. However, it is unclear whether the conventional wisdom of self training would still apply in the case of regurgitative training, because of the difficulty in assessing confidence of LLM outputs (Lin et al.,, 2023). In classification tasks, predicted class probabilities naturally serve as the measure to quantify the uncertainty in a classifier’s predictions. However, LLMs are much more complex than classifiers – they generate multi-token answers in response to prompts. Unless in highly restricted scenarios (e.g., evaluating a single-digit response to the question “what is 2+2”), it is generally not straightforward to define or measure confidence in LLM outputs. As a result, current LLMs do not automatically produce confidence scores for their responses and uncertainty quantification in LLMs remain an open question with many ongoing research, including asking LLMs themselves for “self-evaluation/reflection” (Chen and Mueller,, 2023), re-running the same prompt multiple times and measuring internal consistency among responses (Kotelanski et al.,, 2023), and tapping into human expertise (Shankar et al.,, 2024).
@@ 2.2 Data Augmentation
Data augmentation represents another approach to enrich potentially limited labeled data with synthetic data. It works by injecting noises into existing labeled data instances to artificially create new data instances that can be assigned the same labels. In language tasks, a common data augmentation strategy is back-translation (Yu et al.,, 2018), where a sentence is first translated into a different language and then reverted back to the original language to achieve paraphrasing. In vision tasks, data augmentation may involve image transformation techniques such as rotation, color / contrast modification, etc. (Cubuk et al.,, 2020; Xie et al., 2020a, ). These augmentation strategies can benefit model performance if the injected noises do not change the labels, thereby creating more training data with valid labels.
@@ 2.3 Regurgitative Training
Regurgitative training of generative AI models represents a new problem that has only begun to receive scholarly attention very recently. The earliest work we could identify is Shumailov et al., (2023), which documents that using model-generated data to train next-generation models can create irreversible performance losses, a phenomenon they term “model collapse”. They demonstrate this in common generative AI architectures such as variational autoencoders, Gaussian mixture models, and small-scale LLMs. Moreover, they provide theoretical intuitions that model collapse arises due to errors in model-generated data, which accumulates over more iterations of regurgitative training. Subsequently, the phenomenon of model collapse has also been observed in generative image models (Alemohammad et al.,, 2023; Bertrand et al.,, 2023).
In the meantime, efforts to mitigate model collapse are underway. Bertrand et al., (2023) show that model collapse can be avoided if (i) the proportion of real data is sufficiently high and (ii) model-generated data approximate the distributions of real data well enough. Furthermore, Gerstgrasser et al., (2024) propose to alleviate model collapse by “accumulating data”; that is, using the totality of real and model-generated data (rather than just the model-generated data) to train new models.
We build upon this nascent stream of research and aim to make several distinct contributions. First, we consider regurgitative training of a LLM not only by data generated by itself, but also by other LLMs with varying degrees of capabilities. This is already taking place in practice (e.g., McKinzie et al.,, 2024) but has not been systematically explored in the literature. Second, prior work such as Shumailov et al., (2023) focused on early versions of generative models (e.g., non-transformer-based models or small pre-trained models). Instead, we carry out comprehensive experiments with leading LLMs at the time of research (e.g., GPT-4 and LLAMA2) as well as transformer models trained from scratch, thereby providing a more up-to-date understanding of regurgitative training. Third, we conduct textual analysis to explore the mechanisms that may explain the performance disadvantages of regurgitative training. Finally, we propose several new mitigation strategies beyond what has been tested so far, and empirically evaluate their effectiveness.
@@ 3 Performance Impact of Regurgitative Training
In this section, we aim to understand how regurgitative training affects the performance of an LLM through two sets of experiments, respectively constructed to reflect two representative practices in LLM training: (i) fine-tuning and (ii) training from scratch. Fine-tuning allows users to adapt an existing LLM to their own use cases and, as mentioned before, is a widely adopted practice in the industry (e.g., McKinzie et al.,, 2024). We expect a lot of LLM training will take the form of fine-tuning, because training a state-of-the-art LLM from scratch is highly complex and resource-intensive. Meanwhile, we also consider the case of training smaller-scale transformer language models from scratch, which may be necessary for companies that cannot leverage third-party LLMs due to data security and privacy issues.
@@ 3.1 Experiments with Fine-Tuning
In the first term, $c$ is the length of hypothesis translation and $r$ is the “effective” length of reference translations (defined as the length of the reference translation that best matches the hypothesis translation). This term serves as a “brevity penalty” that assigns a higher score for a better match in lengths between hypothesis and reference translations. In the second term, $p_{n}$ denotes the $n$-gram precision and is defined as
$p_{n}=\frac{\sum_{n\mbox{-}gram\in hypothesis}Count_{matched}(n\mbox{-}gram)}{\sum_{n\mbox{-}gram\in hypothesis}Count(n\mbox{-}gram)}$ (2)
where $Count(n\mbox{-}gram)$ counts the number of $n$-gram in the corpus and $Count_{matched}(n\mbox{-}gram)$ counts the number of $n$-gram matches between the hypothesis translation and reference translations. In Equation (1), the $n$-gram precision scores are then weighted by $w_{n}$ (e.g., uniform weighting $w_{n}=\frac{1}{N}$) to compute the overall BLEU score.
@@
We show the results in Figure 1. Each line represents the translation performance of a particular model over five fine-tuned batches. The $X$-axis indicates batch index, marking the number of data instances utilized in the fine-tuning process. The $Y$-axis represents the BLEU score, where a higher value corresponds to better translation performance.
@@
From the figure, it is evident that the performance of fine-tuning with LLM-generated data (both from the baseline LLM itself and from other LLMs) clearly lags behind the performance of fine-tuning with real human-generated data. Moreover, regurgitative training with different LLMs have differential performance impact. Fine-tuning with data generated by GPT-3.5 itself does not significantly change performance, and fine-tuning with GPT-4 generated data only results in slight performance improvement over the baseline model (i.e., at point 0 on the $X$-axis). However, fine-tuning with LLAMA2 generated data significantly degrades performance compared to the baseline. This is likely because the three LLMs have different translation capabilities. Since we have the ground-truth translations for the 5,000 fine-tuning data, we can directly compute the BLEU scores of translations generated by the three LLMs, and indeed find GPT-4 to be the best ($BLEU=0.3454$), followed by GPT-3.5 ($BLEU=0.3428$) and LLAMA2 ($BLEU=0.2417$).
These results underscore the overall underwhelming, and potentially detrimental effects of regurgitative training. Compared to training with real data, regurgitative training largely stalls learning. Regurgitative training with a better-performing LLM improves performance only marginally and is not sufficient to catch up with the performance on real data. Worse yet, regurgitative training with a less capable LLM can significantly hurt performance.
Note that in the above experiments, we use a small set of data for fine-tuning. This decision stems from the remarkable few-shot learning capabilities of modern LLMs (Brown et al.,, 2020). In addition, we also conduct a robustness check to understand whether the performance of regurgitative training may be different if more fine-tuning data are available. Specifically, we augment the fine-tuning data size by 20 times, to a total of 100,000 data instances, and incrementally add 10,000 per batch. For efficiency and cost considerations, we only run this size-augmented fine-tuning with real data and data generated by GPT-3.5 itself. We then evaluate each fine-tuned models on the same testing data as before. The results are presented in Figure 2. We again observe that regurgitative training is unable to improve translation performance and substantially underperforms training with real data.